OpenStudy (anonymous):

how to prove this trigonometric identity sec x- tan x= 1-sin x/cos x

OpenStudy (anonymous):

i know that sec x equals to 1/cos x so that equation would be 1/cos x -tan x=1-sin x/cos x

OpenStudy (kc_kennylau):

Hints: sec x = 1/cos x tan x = sinx / cosx

OpenStudy (anonymous):

i know that info, but do i subtract the 1/cos x to the other side

OpenStudy (kc_kennylau):

You don't start with an equality to prove an equality

OpenStudy (kc_kennylau):

sec x - tan x = 1/cosx - sinx/cosx = (1-sinx)/cosx

OpenStudy (kc_kennylau):

You can't do this: secx - tanx = (1-sinx)/cosx 1/cosx - tanx = 1/cosx - sinx/cosx 1/cosx - tanx = 1/cosx - tanx

OpenStudy (kc_kennylau):

This is one of the common mistakes that people make when proving anything

OpenStudy (anonymous):

why would the equation become sec x - tan x = 1/cosx - sinx/cosx = (1-sinx)/cosx

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

tan x = a/b = (a/c)/(b/c) = sinx/cosx

OpenStudy (kc_kennylau):

tanx=sinx/cosx will be used often in trigonometry, you'd better memorize it

OpenStudy (anonymous):

after the equation becomes into this sec x - tan x = 1/cosx - sinx/cosx = (1-sinx)/cosx what will we have to do next

OpenStudy (kc_kennylau):

Or more completely: LHS =secx - tanx =1/cosx - sinx/cosx =(1-sinx)/cosx =RHS

OpenStudy (anonymous):

sec -tan x =(1-sinx)/cos x i see it know thank you very much

OpenStudy (kc_kennylau):

no problem